OpenStudy (mrlightbody):

In this question you are asked to determine (x^3)( e^((5x^4)+7)) dx by u substitution

jimthompson5910 (jim_thompson5910):

Hint: let u = 5x^4 + 7 du/dx = 20x^3 du = 20x^3*dx x^3*dx = du/20

OpenStudy (mrlightbody):

Now enter the corresponding integral I in u resulting from your substitution. Your answer should depend only on u

OpenStudy (mrlightbody):

I got what you sent but i dont understand the question I sent.

jimthompson5910 (jim_thompson5910):

hopefully you see how `(x^3)( e^((5x^4)+7)) dx ` turns into `e^u*du/20` ?

OpenStudy (mrlightbody):

yup

OpenStudy (mrlightbody):

here is the real question

OpenStudy (mrlightbody):

@jim_thompson5910 i think you understand more if you see the attachments i posted

jimthompson5910 (jim_thompson5910):

ok let me think

OpenStudy (mrlightbody):

thanks man

jimthompson5910 (jim_thompson5910):

ok for 7.3, I think they just want the transformed integral before you actually do any integrating so you'd have \[\Large \int \frac{1}{20}e^u du\] or \[\Large \frac{1}{20}\int e^u du\]

OpenStudy (mrlightbody):

alright , i will give it a shot

OpenStudy (mrlightbody):

awesome got it right @jim_thompson5910 thanks

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (mrlightbody):

i am stuck on one more question, is it ok that you explain it to me? its in the attachment @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Try \[\Large u = \log_{4}(x)\] so, \[\Large \frac{du}{dx} = \frac{1}{x*\ln(4)}\] \[\Large \ln(4)*du = \frac{dx}{x}\]

jimthompson5910 (jim_thompson5910):

hmm is that +7 inside the log argument? or no?

OpenStudy (mrlightbody):

no it is not. its part of the function @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

so \[\Large \int \frac{(\log_4(x))^2-\sqrt{3}\log_4(x)+7}{x} dx\] \[\Large \int \left((\log_4(x))^2-\sqrt{3}\log_4(x)+7\right)\frac{dx}{x}\] \[\Large \int \left(u^2-\sqrt{3}u+7\right)*\ln(4)*du\] \[\Large \ln(4)\int \left(u^2-\sqrt{3}u+7\right)du\] the second 'u' is not under the square root

OpenStudy (mrlightbody):

thanks, after that I would need to change the integrals to find the value correct ?@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you mean change the limits of integration?

OpenStudy (mrlightbody):

yeah @jim_thompson5910

jimthompson5910 (jim_thompson5910):

if x = 1, then \[\Large u = \log_4(x) = \log_4(1) = ???\]

OpenStudy (mrlightbody):

that equals 0.6021

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

type in `log(1)/log(4)` into your calculator

OpenStudy (mrlightbody):

thats 0.25,

jimthompson5910 (jim_thompson5910):

what calc do you have?

OpenStudy (mrlightbody):

in attachment

jimthompson5910 (jim_thompson5910):

ok hit the log key and hit 1

OpenStudy (mrlightbody):

i get 0

jimthompson5910 (jim_thompson5910):

yes so when you divide by log(4), whatever that is, the overall result of log(1)/log(4) will be 0

jimthompson5910 (jim_thompson5910):

if x = 1, then \[\Large u = \log_4(x) = \log_4(1) = 0\]

jimthompson5910 (jim_thompson5910):

log(64)/log(4) = 3 so if x = 64, then \[\Large u = \log_4(x) = \log_4(64) = 3\]

jimthompson5910 (jim_thompson5910):

\[\large \int_{1}^{64} \frac{(\log_4(x))^2-\sqrt{3}\log_4(x)+7}{x} dx = \ln(4)\int_{0}^{3} \left(u^2-\sqrt{3}u+7\right)du\] take note at how the limits of integration change

OpenStudy (mrlightbody):

ahhh I see.

OpenStudy (mrlightbody):

I will try and solve it from here now and i will send an image you to see if i got it right @jim_thompson5910

OpenStudy (mrlightbody):

is this correct @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

how do they want you to enter the answer? as an exact number? or approximate decimal form?

OpenStudy (mrlightbody):

i got it wrong but they gave me hints.

jimthompson5910 (jim_thompson5910):

it would be nice if they posted the instructions on how to enter the answer. Hmm well if you're using an approximate value, then try 30.7837 that's what I'm getting as an approx value

OpenStudy (mrlightbody):

i think they would like it in exact form. the system uses maple language

jimthompson5910 (jim_thompson5910):

ah ok I see

jimthompson5910 (jim_thompson5910):

in your work, I agree with lines 1 through 3. Line 4 is correct in terms of the right approx answer, but it's best to keep things exact as long as possible so \(\Large \frac{\sqrt{3}(3)^2}{2}\) should turn into \(\Large \frac{9\sqrt{3}}{2}\)

OpenStudy (mrlightbody):

thanks man for the help much! appreciated !! @jim_thompson5910